# Making the OST of the Muon Collider

### v1

@@ IV Physics Case
The OST facility offers a rich physics program that spans a wide range of QCD and BSM phenomena. In contrast to the EIC, which will provide access to the low- regime of nucleon and nuclear structure, OST can deliver high-statistics measurements at higher and moderate-to-large . OST provides unique opportunities to study the valence structure of the nucleon, rigorously utilize perturbative QCD in a regime where non-perturbative corrections are suppressed, and explore new physics with unprecedented sensitivity. In the following subsections we touch briefly upon several potential topics that could be studied at the OST facility.
@@ IV.1 Precision QCD and Parton Distributions
With its high instantaneous luminosity and high beam energy, OST opens the door to a unique regime for measurements of inclusive deep inelastic scattering (DIS). OST will provide ample data for extraction of the valence region parton distribution functions (PDFs) with reduced theoretical uncertainties arising from higher-twist effects and target-mass corrections. Measurements at values up to several thousand GeV2 will provide critical input to global PDF fits, especially in the high- regionĀ BurkertĀ etĀ al. (2023). With a high-power polarized target, the helicity PDFs can be measured with high precision. The high reach of OST minimizes non-perturbative contamination and facilitates clean tests of spin sum rules in QCD.
Furthermore, DIS structure functions have historically provided some of the most powerful constraints on the value of Ā dāEnterriaĀ etĀ al. (2024). Since polarized and nuclear targets can also be leveraged to exploit various QCD sum rules and reduce the typically dominant theoretical uncertaintiesĀ KutzĀ etĀ al. (2024), an extraction of at the level of 1% or better from OST alone seems feasible, to be compared with the global uncertainty of 0.8% from all world measurements combinedĀ dāEnterriaĀ etĀ al. (2024).
@@ IV.3 Exclusive Measurements
Much of the exciting physics at the EIC is driven by multi-dimensional measurements of exclusive processes, which enables āimagingā of hadrons in position space. Since exclusive cross sections usually fall strongly with , OST with a 7 TeV beam is suboptimal. However, at a beam energy of 1.5 TeV, the scattered muon can be measured down to values of to GeV2 at an energy similar to the EIC at x GeV but with almost two orders of magnitude higher luminosity. Measurements of Deeply Virtual Compton Scattering and Deeply Virtual Meson Production at perturbative values of and can complement EIC results at lower- on the spatial and gravitational structure of nucleons and nuclei.
Existing data does not permit the extraction of GPDs, as the measured experimental observables contain combinations of GPDs that cannot yet be uniquely disentangled from one anotherĀ BertoneĀ etĀ al. (2021); MoffatĀ etĀ al. (2023). The field is in need of complementary observables that enable this disentanglement, and beam charge asymmetry measurements are one of the pillars of the CEBAF positron upgrade for this reasonĀ VoutierĀ etĀ al. (2023); ZhaoĀ etĀ al. (2021); AccardiĀ etĀ al. (2021). If the Muon Collider complex is constructed such that both muons and antimuons can be delivered to the same experiment, multi-dimensional beam charge asymmetries of exclusive processes such as double-DVCS can be measured at perturbative scales to help solve the problem of extracting GPDsĀ MoffatĀ etĀ al. (2023). OST is particularly well-suited for double-DVCS by virtue of the high luminosity and ability to precisely measure either muons or electrons from the decay of the virtual photon.
@@ IV.5 BSM Searches
A further exciting opportunity offered by OST is the ability to perform high-precision tests of the Standard Model in an unexplored kinematic regime. With billions of DIS events at GeV2, OST can perform stringent tests of lepton universality, comparing muon-induced scattering to electron-induced processes. The high regime is especially sensitive to contact interactions that could arise from new physics beyond the Standard Model. In the language of Standard Model Effective Field Theory (SMEFT), precise measurements of differential cross sections and asymmetries can constrain higher-dimensional operators sensitive to new physicsĀ BoughezalĀ etĀ al. (2020); BissolottiĀ etĀ al. (2023); BoughezalĀ etĀ al. (2022); CiriglianoĀ etĀ al. (2021). This sensitivity is complementary to that provided by collider experiments, offering independent tests of BSM scenarios.
Finally, OSTās kinematics provide an ideal environment to study electroweak interference effects in DISĀ Pro (2012); AbramowiczĀ etĀ al. (2015); AndreevĀ etĀ al. (2018), which can yield precision determinations of the weak mixing angle at momentum transfers far larger than existing measurements in lepton scattering.

### v2

@@ IV Physics Case
The OST facility offers a rich physics program that spans a wide range of QCD and BSM phenomena. In contrast to the EIC, which will provide access to the low- regime of nucleon and nuclear structure, OST can deliver high-statistics measurements at higher and moderate-to-large as shown in Figs. 3. OST provides unique opportunities to study the valence structure of the nucleon, rigorously utilize perturbative QCD in a regime where non-perturbative corrections are suppressed, and explore new physics with unprecedented sensitivity. In the following subsections we touch briefly upon several potential topics that could be studied at the OST facility.
@@ IV.1 Precision QCD and Parton Distributions
With its high instantaneous luminosity and high beam energy, OST opens the door to a unique regime for measurements of inclusive deep inelastic scattering (DIS). OST will provide ample data for extraction of the valence region parton distribution functions (PDFs) with reduced theoretical uncertainties arising from higher-twist effects and target-mass corrections. Measurements at values up to several thousand GeV2 will provide critical input to global PDF fits, especially in the high- region Burkert:2022hjz. With a high-power polarized target, the helicity PDFs can be measured with high precision. The high reach of OST minimizes non-perturbative contamination and facilitates clean tests of spin sum rules in QCD.
Furthermore, DIS structure functions have historically provided some of the most powerful constraints on the value of dEnterria:2022hzv. Since polarized and nuclear targets can also be leveraged to exploit various QCD sum rules and reduce the typically dominant theoretical uncertainties Kutz:2024eaq, an extraction of at the level of 1% or better from OST alone seems feasible, to be compared with the global uncertainty of 0.8% from all world measurements combined dEnterria:2022hzv.
@@ IV.3 Exclusive Measurements
Much of the exciting physics at the EIC is driven by multi-dimensional measurements of exclusive processes, which enables ”imaging” of hadrons in position space. Since exclusive cross sections usually fall strongly with , OST with a 7 TeV beam is suboptimal. However, at a beam energy of 1.5 TeV, the scattered muon can be measured down to values of to GeV2 at an energy similar to the EIC at x GeV but with almost two orders of magnitude higher luminosity. Measurements of Deeply Virtual Compton Scattering and Deeply Virtual Meson Production at perturbative values of and can complement EIC results at lower- on the spatial and gravitational structure of nucleons and nuclei.
Existing data does not permit the extraction of GPDs, as the measured experimental observables contain combinations of GPDs that cannot yet be uniquely disentangled from one another Bertone:2021yyz; Moffat:2023svr. The field is in need of complementary observables that enable this disentanglement, and beam charge asymmetry measurements are one of the pillars of the CEBAF positron upgrade for this reason CLAS:2023gja; Zhao:2021zsm; Accardi:2020swt. If the Muon Collider complex is constructed such that both muons and antimuons can be delivered to the same experiment, multi-dimensional beam charge asymmetries of exclusive processes such as double-DVCS can be measured at perturbative scales to help solve the problem of extracting GPDs Moffat:2023svr. OST is particularly well-suited for double-DVCS by virtue of the high luminosity and ability to precisely measure either muons or electrons from the decay of the virtual photon.
@@ IV.5 BSM Searches
A further exciting opportunity offered by OST is the ability to perform high-precision tests of the Standard Model in an unexplored kinematic regime. With billions of DIS events at GeV2, OST can perform stringent tests of lepton universality, comparing muon-induced scattering to electron-induced processes. The high regime is especially sensitive to contact interactions that could arise from new physics beyond the Standard Model. In the language of Standard Model Effective Field Theory (SMEFT), precise measurements of differential cross sections and asymmetries can constrain higher-dimensional operators sensitive to new physics Boughezal:2020uwq; Bissolotti:2023vdw; Boughezal:2022pmb; Cirigliano:2021img. This sensitivity is complementary to that provided by collider experiments, offering independent tests of BSM scenarios.
Finally, OST’s kinematics provide an ideal environment to study electroweak interference effects in DIS Proceedings:2012ulb; H1:2015ubc; H1:2018mkk, which can yield precision determinations of the weak mixing angle at momentum transfers far larger than existing measurements in lepton scattering.